AA borrower takes out a 29-year mortgage loan for $286,819 with an interest rate of 9%. What would the monthly payment be?

BA borrower takes out a 30-year mortgage loan for $190,372 with an interest rate of 8% and monthly payments. What portion of the first months payment would be applied to interest?

CA borrower has a 25-year mortgage loan for $495,186 with an interest rate of 9% and monthly payments. If she wants to pay off the loan after 8 years, what would be the outstanding balance on the loan?
Answer #1
A) Monthly payment (using formula for loan amortization) = 286819*(0.09/12)*(1+0.09/12)^348/((1+0.09/12)^348-1)) = $         2,323.69
B) Monthly payment (using formula for loan amortization) = 190372*(0.08/12)*(1+0.08/12)^360/((1+0.08/12)^360-1)) = $         1,396.88
Interest on the first installment = 190372*0.08/12 = $         1,269.15
Portion of installment = 1269.15/1396.88 = 90.86%
C) Monthly payment (using formula for loan amortization) = 495186*(0.09/12)*(1+0.09/12)^300/((1+0.09/12)^300-1)) = $         4,155.58
Outstanding balance after 8 years is the PV of the remaining installments = 4155.58*((1+0.09/12)^204-1))/((0.09/12)*(1+0.09/12)^204)) = $   4,33,412.92